Go from 0 to 1 on the x-axis, then back half-way to 1/2, then forward half as far to 3/4, then back as half as far to 5/8, then forward half as far, and so on. Give an explicit formula for your position after the nth step. Where do you end up?


1
Expert's answer
2021-01-29T00:08:33-0500

1) An = 0 + 20 - 2-1 + 2-2 - 2-3 + ... + 2-(n-1) , where n = "2i-1", "i\in\N" (for odd).

Because at the odd step we move forward, so "plus".

An = (20 - 2-1) + (2-2 - 2-3) + ... + (2-(n-3) - 2-(n-2)) + 2-(n-1)

An = (2*2-1 - 2-1) + (2*2-3 - 2-3) + ... + (2*2-(n-2) - 2-(n-2)) + 2-(n-1)

An = (2-1) + (2-3) + ... + (2-(n-2)) + 2-(n-1)


2) An = 0 + 20 - 2-1 + 2-2 - 2-3 + ... - 2-(n-1) , where n = "2i", "i\in\N" (for even).

Because at an even step we move backward, so "minus".

An = (20 - 2-1) + (2-2 - 2-3) + ... + (2-(n-4)) - 2-(n-3)) + (2-(n-2)) - 2-(n-1))

An = (2*2-1 - 2-1) + (2*2-3 - 2-3) + ...+ (2*2-(n-3)) - 2-(n-3)) + (2*2-(n-1) - 2-(n-1))

An = (2-1) + (2-3) + ...+ (2-(n-3)) + (2-(n-1))


If it is necessary to bring to a common denominator, then :

1)

An "=\frac{2^{-1}*2^{n-1}+2^{-3}*2^{n-1}+...+2^{-(n-2)}*2^{n-1}+1}{2^{n-1}}"

An "=\frac{2^{n-1-1}+2^{n-1-3}+...+2^{n-1-(n-2)}+1}{2^{n-1}}"

An "=\frac{2^{n-2}+2^{n-4}+...+2^1+1}{2^{n-1}}" , for odd.

2)

An "=\frac{2^{-1}*2^{n-1}+2^{-3}*2^{n-1}+...+2^{-(n-3)}*2^{n-1}+1}{2^{n-1}}"

An "=\frac{2^{n-2}+2^{n-4}+...+2^2+1}{2^{n-1}}" , for even.